1. A billiard ball moves 18 meters southward, then 12 meters northward, and finally 8 meters southward.

For this motion, what is the distance moved? _________________

What is the magnitude and direction of the displacement? _________________

2. A horse gallops 36 meters eastward, then 44 meters westward, and finally 22 meters eastward.

For this motion, what is the distance moved? _________________

What is the magnitude and direction of the displacement? _________________

3. A car travels 20 meters north, 15 meters south and finally 45 meters north

For this motion, what is the distance moved? _________________

What is the magnitude and direction of the displacement? _________________


1
Expert's answer
2021-11-14T17:17:56-0500

1. A billiard ball moves 18 meters southward, then 12 meters northward, and finally 8 meters southward.

For this motion, what is the distance moved? 18+12+8=38 m

What is the magnitude and direction of the displacement? 18-12+8=14 m southward

2. A horse gallops 36 meters eastward, then 44 meters westward, and finally 22 meters eastward.

For this motion, what is the distance moved? 36+44+22=102 m

What is the magnitude and direction of the displacement? 36-44+22=14 m eastward


3. A car travels 20 meters north, 15 meters south and finally 45 meters north

For this motion, what is the distance moved? 20+15+45=80 m

What is the magnitude and direction of the displacement? 20-15+45=50 m northward